Willie Company sells 32,000 units at $25 per unit. Variable costs are $14.25 per unit, and fixed costs are $175,400. Determine (a) the contribution margin ratio, (b) the unit contribution margin, and (c) income from operations. a. Contribution margin ratio (Enter as a whole number.) % b. Unit contribution margin (Round to the nearest cent.) $ per unit c. Income from operations
Answer #1

Contribution margin=Sales-Variable cost  

=(25-14.25)=$10.75 per unit

Contribution margin ratio=Contribution margin/Sales

=10.75/25

=43%

Unit contribution margin=$10.75 per unit

Total contribution margin=(10.75*32000)=$344000

Less:Fixed costs=(175400)

Income from operations=$168600